A diffraction grating has 6.9×105 slits/m.

Q: Find the angular spread in the second-order spectrum between red light of wavelength 6.8×10−7 m and blue light of wavelength 4.3×10−7 m .

Express your answer to two significant figures and include the appropriate units.
Answer #1

when applied force is slightly greater than 3N the book set in motion.

width of each slit,

d = 1/(6.9*10^5)

= 1.449*10^-6 m

second order for red light

d*sin(theta_red) = 2*lamda_red

sin(theta_red) = 2*lamda_red/d

= 2*6.8*10^-7/(1.449*10^-6)

= 0.9385783

theta_red = sin^-1(0.9385783)

= 69.814 degrees

second order for blue light

d*sin(theta_blue) = 2*lamda_blue

sin(theta_blue) = 2*lamda_blue/d

= 2*4.3*10^-7/(1.449*10^-6)

= 0.59351

theta_blue = sin^-1(0.59351)

= 36.406 degrees

the angular spread in the second-order spectrum between red and blue light,

delta_theta = theta_red - theta_blue

= 69.814 - 36.406

= 33.4 degrees

= 33 degrees <<<<<<<<<<------------------------------Answer